9) BankOne has reserves of $120, loans at $380 and checkable deposits of $500. The required...

9)

BankOne has reserves of $120, loans at $380 and checkable deposits of $500. The required reserve ratio

is 12%. A customer withdraws $40 from her account. After the withdrawal, the bank has excess

reserves of _____.

10)

Suppose a market for financial assets has type Aa and type Bb. Buyers value Aa at $14,500 and type Bb

at $11,500, while sellers value Aa at $12,700 and Bb at $10,100. If the buyers cannot observe type, what

is the min fraction of type As needed in the market to avoid adverse selection?
Answer #1

9) Required reserve ratio = 12%

Required reserves = 12% of 500 = 60

Excess reserves = resrves - required reserves = 120-60 = 60

After the withdrawal of $40, Deposits = 500-40 =460

Required reserves = 12% of 460 = 55.2

Excess reserves = 120-55.2 = 64.8

As per Chegg guidelines, only the first question can be answered.